# TaskEval: Assessing Difficulty of Code Generation Tasks for Large Language Models

### v2

@@ 5.3. RQ3: Do programming constructs affect tasks’ characteristics?
Results of the correlations between node types extracted and tasks’ characteristics as well as AD test are given in Table 1 and Table 2. To facilitate interpretation, we regrouped node types into semantically similar categories before calculating the categories. Hence, Op includes all nodes related to math operations such as addition, subtraction or even boolean operations, Loop includes both For and While node types, DataStruct contains all List, Set, Dict or Tuple node types etc. All details on the grouping process are available in our replication package (rep, 2024).
@@
Observing the results, we distinguish several patterns that can be intra-dataset (across CodeLLMs) or inter-dataset (between HumanEval+ and ClassEval). First, looking at the comparison with tasks’ difficulty (Table 1), several node types such as If and Op are present across CodeLLMs and benchmarks, highlighting a commonality in the programming constructs that CodeLLMs use to address harder/easier tasks. While harder tasks could result in longer code which might bias the outcome, as we controlled for the size of the code fragments, it points out that it is not simply about more absolute processing but that to tackle harder tasks CodeLLMs rely on a higher level of processing. As such, the harder the tasks, the more the CodeLLMs need conditions and variable assignments/operations, independently of the number of lines of code to solve the task. In turn, this might increase their chance of producing erroneous code as they get lost in the logic flow. Note that this highlights how CodeLLMs attempt to solve the task, rather than what is required to solve the task. There are various ways to approach each task, and optimizing the code was not a requirement for CodeLLMs. As a result, programming constructs such as If, Op, Assign, and Loop show positive correlations, along with significant statistical differences between the distributions of programming constructs for the top 50% easiest and hardest tasks, despite fluctuations in the strength of these correlations and significance. In contrast, some differences emerge between the two benchmarks, highlighted by stronger correlations and/or more significant results from the AD test for CodeLLMs. For HumanEval+, CodeLLMs tend to utilize more FunctionDef (defining additional sub-functions) and Comparison (e.g., Compare) constructs when tackling harder tasks. For ClassEval, CodeLLMs make more use of index manipulation (Subscript) or structure comprehension (comprehension such as list comprehension, and dict comprehension). In that last case, we find non-significant correlations for this same node type for HumanEval+, despite said node type being used in HumanEval+ tasks. In that case, it means that this is not a property of tasks in general but it is a property of the tasks of ClassEval. Correlations are mostly weak ($>0.06$), with only the correlations with the If node in HumanEval+ being moderate ($>0.26$) (Gilpin, 1993; Schober et al., 2018). It should be noted that it is possible to have, for most CodeLLMs, significant AD test without significant correlations (e.g., Subscript in HumanEval+) or the invert (e.g., DataStruct in ClassEval): Certain node types may be more prevalent in harder tasks, but since the AD test measures the difference between the top and bottom 50%, and the median is relatively low, this can result in a nonsignificant difference.
Regarding CodeLLMs, we observe that GPT-3.5 exhibits generally fewer significant correlations between task difficulty and the programming constructs used. On the contrary, correlations seem more numerous overall for Code Llama and Code Gemma. Moreover, in terms of AD test, GPT3.5 has the lowest number of statistical differences between $50\%$ easiest/hardest tasks in HumanEval+ and is tied in ClassEval. Moreover, we observe that for some node types, some CodeLLMs do or do not exhibit a correlations/AD difference while the other CodeLLMs exhibit the opposite. For instance, in HumanEval+, Raise and Subscript nodes exhibit a positive correlation for all CodeLLMs except GPT3.5. This highlights that CodeLLMs will deal with tasks within the benchmark differently, based on their understanding of the prompt, their training dataset, and their architecture. While this can lead to errors, as we saw for instance in Listing 6 in the previous Section 5.2, it can also be harmless but emphasize particularity in how the prompts are processed. For instance, for the Raise node type in HumanEval+, we note that they are most often used by CodeLLMs to handle constraints in the prompts that are informative (e.g., “the numbers will be between 0 and 1000”). These constraints are not enforced in the tests (and so with or without, the code is correct) but are phrased in the prompts which will be reflected in how the CodeLLMs will generate a code. In this case, for several tasks, GPT3.5 will return a message as a string rather than raising an exception. The correlation tends to be positive as tasks with these informative constraints tend to be harder.
@@
Benchmark Node Deepseek MagiCoder GPT3.5 CodeGemma CodeLLama
$\tau$ AD $\tau$ AD $\tau$ AD $\tau$ AD $\tau$ AD
@@
[…]
Compare 0.112∗∗ -- 0.133∗∗ $\ast$ 0.103 -- 0.138∗∗∗ -- 0.103 --
[…]
Regarding now the comparison with the tasks’ discriminant, we note that they are no cross CodeLLMs’ patterns for HumanEval+ and so tasks’ discriminant does not seem to influence the type of program constructs used. However, for ClassEval, we note a few intra-benchmarks patterns. First, both AD and correlations are significant for all CodeLLMs when it comes to Assign node type. Given that the correlations are negative, this points towards CodeLLMs using variable assignments less on high discriminant tasks compared to low discriminant tasks (i.e., where CodeLLMs have a similar probability of being correct). This might be a contributing factor to why such tasks are less discriminant. That is, the higher number of assignments needed (or at least used by CodeLLMs) can increase the logic and so the likelihood of the CodeLLMs making a mistake, with more capable CodeLLMs being less likely to make a mistake compared to less capable ones as per the high discriminant. We also found that most CodeLLMs feature significant correlations/AD on comprehension and DataStruct node types with GPT3.5 being the only one not featuring such behavior in both cases.
[…]
CodeLLM Node Deepseek MagiCoder GPT3.5 CodeGemma CodeLLama
$\tau$ AD $\tau$ AD $\tau$ AD $\tau$ AD $\tau$ AD
HumanEval+ Lambda - -- - -- 0.119 -- 0.117 -- 0.122 --
Op - -- - -- -0.104 -- - -- - --
[…]
Attribute - -- - -- - -- 0.102 -- - --
[…]
Compare -0.093 -- -0.092 -- - -- - -- - --
Try-Catch - -- - -- 0.109 -- 0.125∗∗ -- 0.111 --
Import - -- - -- - -- - -- -0.122 --
Findings 3: Cross-checking AST of generated codes with tasks’ characteristics points towards specific program constructs used by CodeLLMs to address a task. Conditions, variables assignments, and loop are generally used more in harder tasks, independently of the code length, and some node types are more specific to a particular benchmark (e.g., Subscript in ClassEval and Compare in HumanEval+). We do not find any particular patterns when using the discriminant of the tasks for both benchmarks.
@@ 5.4. RQ4: Do human annotators and CodeLLMs rate tasks’ difficulty similarly?
As detailed in Section 4.5, estimating difficulty of tasks has been conducted in several iterative rounds. After the first round, only $6$ tasks did not reach the aimed agreement. In the second round, we thus redid one questionnaire with these $6$ tasks, following the same template, except we added the minimum/median/maximum time/difficulty estimations of the previous round, as well as comments provided by previous participants. We obtained again $5$ answers and computed the agreement resulting in only $1$ tasks not reaching the threshold. Analysis of this task revealed that all but one participant judged the task long (40 minutes or more) and one participant judged it quick (5 minutes), which resulted in an agreement below $0.6$ because of the metric used in Alhamed et al. (Alhamed and Storer, 2021), the same as the first round. Given this observation, even with a lower agreement, we accepted the task as most participants agreed on the time effort.
[…]
@@
We then compared the human annotators’ difficulty assessment and the CodeLLMs’ assessment. Given that we have two types of rating on the same items, we can use the Cohen-$\kappa$ to measure the agreement between the two raters (i.e., the human annotators median and the IRT based difficulty). The agreement is weighted using a linear scheme that is, for instance, rating a task as 00 and $1$ will be considered as a better agreement than 00 and $2$. This results in a $\kappa=0.21$ for HumanEval+ and a $\kappa=0.14$ for ClassEval, resulting in respectively a Slight and Fair agreement (Munoz and Bangdiwala, 1997), which is a relatively low agreement. As such this points toward human annotators not being a reliable proxy for the difficulty as estimated by CodeLLMs, as there exists a non-trivial number of tasks on which both sides disagree. While Alhamed et al. (Alhamed and Storer, 2021) showed that annotators tend to underestimate the real difficulty of the tasks, we found that this does not affect our result. Indeed, even if we were to increase the difficulty reported by human annotators by one category (i.e., accounting for the human annotators underestimating the real difficulty), the $\kappa$ coefficient of agreement between the human annotators and the CodeLLMs’ assessment would remain unchanged. We found a similar results if we consider that human annotators overestimate the difficulty.
@@
Table 3. Examples of difficulty ratings between humans and LLMs for some tasks of HumanEval+ (Left) and ClassEval (Right). Numbers represent the reported difficulty on the scale used by our human assessors. We discretize the difficulty evaluated for our LLMs on the same scale for comparison.
@@
Encoding by shifting characters 4 2 Hexadecimal conversion 0 2
Prime factors of an integer 1 3 XML files write 0 3
Triplet under conditions 2 4 SQL requests on age 4 1
List with sentences as elements 5 2 Convert Unicodes in HTML 5 3
Thus, this shows that the estimation by human annotators can, at least in several cases, drastically differ from CodeLLMs. For instance, in HumanEval+, several of the easiest tasks for CodeLLMs that were assessed as harder by participants are mathematical/logical in nature, for example sorting tasks based on a condition (like binary representation of the numbers and prime numbers). In this case, while the tasks themselves are not inherently difficult, they become more complex due to additional concepts that human annotators perceive as increasing the difficulty. This might, of course, differ depending on the annotators, yet the ratings we aim to capture is a general assessment. In contrast, tasks with convoluted prompts or multiple conditions, which were challenging for CodeLLMs, tended to be rated simpler by humans. We give examples of such tasks in Table 3.
In that case, CodeLLMs might struggle to keep track of all conditions. We give an example in Listing 7. In that example, CodeLLMs’ difficulty is high (0.96) while human-annotators’ median announced time is 5 min (low difficulty). Human-annotators point out that the task is easy, as it is a basic string manipulation with a simple edge case. Observing CodeLLMs’ codes, it is clear that the CodeLLMs struggle handling the two conditions, especially the last one. Indeed, the CodeLLMs’ code generally messes the two conditions up by merging them together (e.g., “Example 1” should be “Example-1”, but the generated code generally returns “Example-__1”).
@@
Listing 7: Example of a task, represented here by a single prompt, difficult for CodeLLMs but judged simple by human-annotators.
Findings 4: Human-annotators assessment differ in both case to CodeLLMs’ difficulty assessment, emphasizing that using human judgement to assess CodeLLMs’ tasks difficulty might not be a valid proxy. This motivates the usage of approach like ours that rely directly on CodeLLMs’ assessments to establish characteristics of the tasks.

### v3

@@ 5.4. RQ3: Do programming constructs affect tasks’ characteristics?
Results of the correlations between node types extracted and tasks’ characteristics, as well as AD test are given in Table 1 and Table 3 for HumanEval+, and in Table 2 for ClassEval (we did not find any significant correlation between discriminants and program constructs for ClassEval). To facilitate interpretation, we regrouped node types into semantically similar categories before calculating the categories. Hence, Op includes all nodes related to math operations such as addition, subtraction or even boolean operations, Loop includes both For and While node types, DataStruct contains all List, Set, Dict or Tuple node types, etc. All details on the grouping process are available in our replication package (rep, 2024).
@@
Observing the results, we distinguish several patterns that can be intra-dataset (across CodeLLMs) or inter-dataset (between HumanEval+ and ClassEval). First, looking at the comparison with tasks’ difficulty (Table 1 and Table 2), several node types such as If and Assign are present across CodeLLMs and benchmarks, highlighting a commonality in the programming constructs that CodeLLMs use to address harder/easier tasks. While harder tasks could result in longer code, which might bias the outcome, as we controlled for the size of the code fragments, it points out that it is not simply about more absolute processing, but that to tackle harder task,s CodeLLMs rely on a higher level of processing. Moreover, since the trend is observed at different similarity thresholds, this points towards an intrinsic property of easy/hard tasks. As such, the harder the tasks, the more the CodeLLMs need conditions (especially comparison operators, given the widespread significance of Compare in both benchmarks) independently of the number of lines of code to solve the task. Additionally, increasingly harder tasks require more manipulation of data through variable assignment. In turn, this might increase their chance of producing erroneous code as they get lost in the logic flow. Note that this highlights how CodeLLMs attempt to solve the task, rather than what is required to solve the task. There are various ways to approach each task, and optimising the code was not a requirement for CodeLLMs.
In contrast, some differences emerge between the two benchmarks, highlighted by stronger correlations and/or more significant results from the AD test for CodeLLMs. For HumanEval+, CodeLLMs tend to utilise more DataStruct (e.g., list, dict etc.) constructs when tackling harder tasks. For ClassEval, CodeLLMs make more use of index manipulation (Subscript) or operators (Op such as boolean ’or’, binary ’not’ or simple addition ’+’). In that last case, we find non-significant correlations for this same node type for HumanEval+, despite said node type being used in HumanEval+ tasks. In that case, it means that this is not a property of tasks in general, but it is a property of the tasks of ClassEval. Correlations are mostly weak (>0.06>0.06), with only the correlations with the If node in HumanEval+/ClassEval and Op for ClassEval being moderate (>0.26>0.26) (Gilpin, 1993; Schober et al., 2018). It should be noted that it is possible to have significant AD test without significant correlations (e.g., Lambda in HumanEval+) or the inverse (e.g., arg in ClassEval). This can arise when there are differences between the distribution of the top/bottom 50% tasks’ difficulty, but without a clear trend in terms of difficulty, and so no correlations.
Regarding CodeLLMs, we observe that for some node types, some CodeLLMs do or do not exhibit a correlation/AD difference, while the other CodeLLMs exhibit the opposite. For instance, in HumanEval+, FunctionDef nodes exhibit a correlation for all but DeepSeek-V3 and QwenCoder, highlighting that these models might not rely on ancillary functions to solve HumanEval’s harder tasks. In ClassEval, DeepSeek-V3 is the only CodeLLM without a significant correlation/AD for comprehension (e.g. list comprehension, etc.). This highlights that CodeLLMs will deal with tasks within the benchmark differently, based on their understanding of the prompt, their training dataset, and their architecture. While this can lead to errors, as we saw for instance in Listing 7 in the previous Section 5.3, it can also be harmless but emphasise particularity in how the prompts are processed.
@@
Node CodeLLMs
DeepSeekCoder MagiCoder CodeLlama GPT3.5 CodeGemma DeepSeek-V3 QwenCoder YiCoder
$\tau$ AD $\tau$ AD $\tau$ AD $\tau$ AD $\tau$ AD $\tau$ AD $\tau$ AD $\tau$ AD
Assign .24/.26 * .22/.25 * .25/.27 * .22/.24 * .18/.22 * .26/.28 * .22/.24 * .23/.25 *
Lambda - * - * - - - * - - - - - - - -
Subscript .14/.16 * .15/.17 * .14/.16 * - - .14/.16 - - - .13/.15 - .17/.18 *
If .27/.30 * .26/.29 * .24/.27 * .27/.31 * .25/.29 * .28/.30 * .26/.30 * .26/.31 *
Loop .12/.14 - .13/.16 * .13/.16 - .13/.16 * - - .13/.15 - .14/.15 * .12/.16 -
Compare .14/.16 * .12/.15 * - - .13/.16 * .11/.14 * .10/.12 * .13/.14 * .14/.15 *
comprehension - * - - - - - - - * - - - - - -
FunctionDef .16/.25 * .13/.21 * .13/.27 * .15/.20 * .13/.23 * - - - - .13/.24 *
DataStruct .15/.19 * .14/.18 * .17/.21 * .14/.17 * .15/.19 * .17/.18 * .13/.14 * .13/.17 *
@@
Node CodeLLMs
DeepSeekCoder MagiCoder CodeLlama GPT3.5 CodeGemma DeepSeek-V3 QwenCoder YiCoder
$\tau$ AD $\tau$ AD $\tau$ AD $\tau$ AD $\tau$ AD $\tau$ AD $\tau$ AD $\tau$ AD
arg .14/.16 * .11/.14 - - - .12/.15 - .11/.15 - .12/.14 - .13/.16 - .12/.13 -
Assign .14/.24 - .17/.27 * .17/.26 * .19/.29 * .12/.24 - .16/.25 * .14/.25 * .18/.29 *
Subscript .20/.31 * .19/.28 * .19/.29 * .18/.27 * .21/.29 * .17/.25 * .20/.27 * .19/.31 *
If .25/.28 * .25/.26 * .24/.27 * .25/.27 * .25/.27 * .28/.29 * .27/.29 * .28/.31 *
Loop - - .16/.24 * .12/.25 - .14/.23 - - - .14/.23 * .14/.24 * .13/.28 -
Attribute - - - - - - - - - - - * - - - -
Compare .17/.22 * .17/.21 * .16/.18 * .16/.20 * .19/.22 * .19/.21 * .16/.19 * .20/.24 *
comprehension .17/.27 * .14/.26 - .15/.24 - .16/.24 - .15/.22 * - - .14/.24 * .12/.24 -
DataStruct .12/.23 * .12/.22 * .11/.20 * - - - - - - .14/.24 * .16/.27 *
Op .25/.30 * .23/.29 * .29/.31 * .25/.30 * .28/.31 * .24/.28 * .28/.31 * .28/.33 *
Regarding the comparison with the tasks’ discriminant, we note that there are a few cross-CodeLLMs’ patterns for HumanEval+, namely Assign, Attributes and DataStruct. Both Assign and DataStruct were already significant for tasks’ difficulty. We might explain this behaviour by looking at Figure 4: as the cut-off for top/bottom 50% most/least discriminant tasks is around 1, we see that highly discriminant tasks tend to be, on average, more difficult than low-discriminant tasks for HumanEval+. As such, we might observe a similar pattern in the case of HumanEval+ because of this phenomenon. In the case of HumanEval+, the more discriminant the task, the more likely variable assignment and data structure will be used. At the same time, the more code samples will contain attribute usage (e.g. “mylist.append()” where append is an attribute of the list). On the other hand, we do not observe any patterns for ClassEval.
@@
Node CodeLLMs
DeepSeekCoder MagiCoder CodeLlama GPT3.5 CodeGemma DeepSeek-V3 QwenCoder YiCoder
$\tau$ AD $\tau$ AD $\tau$ AD $\tau$ AD $\tau$ AD $\tau$ AD $\tau$ AD $\tau$ AD
Assign .15/.19 * .15/.18 * .15/.20 * .13/.17 * .14/.19 * .13/.18 * .17/.20 * .15/.19 *
Lambda - - - * - - - * - - - - - - - -
Subscript - * - * - * - * - * - * - * - *
Attribute .15/.18 * .15/.18 * .18/.20 * .11/.13 * .14/.15 * .14/.17 * .17/.19 * .15/.16 *
Compare - * - * - - - - - - - * - - - -
comprehension - - - - - - - - - * - - - - - -
Call - - - - - - - - - - - * - - - -
FunctionDef - - .14/.19 * - * .18/.24 * - - - - .13/.18 * - -
DataStruct .11/.15 * .13/.15 * .14/.18 * .13/.15 * .16/.18 * .15/.17 * .11/.13 * .14/.16 *
Op - * - * - - - * - - - * - * - -
Findings 3: Cross-checking the AST of generated codes with tasks’ characteristics points towards specific program constructs used by CodeLLMs to address a task. Conditions and variable assignment are generally used more in harder tasks, independently of the code length, while some node types are more specific to a particular benchmark (e.g., Subscript in ClassEval and DataStruct in HumanEval+). We do not find any particular patterns when using the discriminant of the tasks for both benchmarks, though we observe correlations for the same specific node types for HumanEval+ (i.e. Assign, Attribute, etc.).
@@ 5.5. RQ4: Do human annotators and CodeLLMs rate tasks’ difficulty similarly?
As detailed in Section 4.5, estimating the difficulty of tasks has been conducted in several iterative rounds. After the first round, only $6$ tasks did not reach the aimed agreement. In the second round, we thus redid one questionnaire with these $6$ tasks, following the same template, except we added the minimum/median/maximum time/difficulty estimations of the previous round, as well as comments provided by previous participants. We obtained again $5$ answers and computed the agreement, resulting in only $1$ task not reaching the threshold. Analysis of this task revealed that all but one participant judged the task long (40 minutes or more), and one participant judged it quick (5 minutes), which resulted in an agreement below $0.6$ because of the metric used in Alhamed et al. (Alhamed and Storer, 2021), the same as the first round. Given this observation, even with a lower agreement, we accepted the task as most participants agreed on the time effort.
To analyse the results, we first compare the human annotators’ difficulty assessment between HumanEval+ and ClassEval. The AD test between both distributions reveals a pvalue<0.01-value<0.01, showing a significant difference between the two distributions. When examining the distributions, we found that annotators tended to rate tasks from ClassEval more difficult than HumanEval+: $40\%$ of ClassEval were labelled as taking $20$ minutes or more, while only $17\%$ of HumanEval+ were labelled as requiring the same time. This contrasts with the assessment made by CodeLLMs, as we saw in RQ1 and RQ,2 that HumanEval+ tended to have a higher number of medium-difficult tasks for the CodeLLMs compared to ClassEval. While this could be attributed to a bias in our sampled tasks, we found no significant difference between the difficulty distributions of our sampled tasks and those of the entire set of tasks.
@@
We then compared the human annotators’ difficulty assessment and the CodeLLMs’ assessment. Given that we have two types of rating on the same items, we can use the Cohen-$\kappa$ to measure the agreement between the two raters (i.e., the human annotators’ median and the IRT-based difficulty). The agreement is weighted using a linear scheme, that is, for instance, rating a task as 0 and $1$ will be considered as a better agreement than 0 and $2$. This results in a $\kappa=0.21$ for HumanEval+ and a $\kappa=0.17$ for ClassEval, resulting in low agreements, respectively a Slight and Fair agreement (Munoz and Bangdiwala, 1997). As such, this points toward human annotators not being a reliable proxy for the difficulty as estimated by CodeLLMs, as there exists a non-trivial number of tasks on which both sides disagree. While Alhamed et al. (Alhamed and Storer, 2021) showed that annotators tend to underestimate the real difficulty of the tasks, we found that this does not affect our result. Indeed, even if we were to increase the difficulty reported by human annotators by one category (i.e., accounting for the human annotators underestimating the real difficulty), the $\kappa$ coefficient of agreement between the human annotators and the CodeLLMs’ assessment would not change significantly. We found similar results if we consider that human annotators overestimate the difficulty.
@@
Table 4. Examples of difficulty ratings between humans and LLMs for some tasks of HumanEval+ (Left) and ClassEval (Right). Numbers represent the reported difficulty on the scale used by our human assessors. We discretise the difficulty evaluated for our LLMs on the same scale for comparison.
@@
Encoding by shifting characters 5 2 Hexadecimal conversion 0 2
Prime factors of an integer 0 3 XML files write 1 3
Triplet under conditions 2 4 Parse and split url 4 2
List with sentences as elements 1 2 Convert Unicodes in HTML 5 3
Thus, this shows that the estimation by human annotators can, at least in several cases, drastically differ from CodeLLMs. For instance, in HumanEval+, several of the easiest tasks for CodeLLMs that were assessed as harder by participants are mathematical/logical in nature, for example, sorting tasks based on a condition (like binary representation of the numbers and prime numbers). In this case, while the tasks themselves are not inherently difficult, they become more complex due to additional concepts that human annotators perceive as increasing the difficulty. This might, of course, differ depending on the annotators, yet the ratings we aim to capture are a general assessment. In contrast, tasks with convoluted prompts or multiple conditions, which were challenging for CodeLLMs, tended to be rated simpler by humans. We give examples of such tasks in Table 4.
In that case, CodeLLMs might struggle to keep track of all conditions. We give an example in Listing 8. In that example, CodeLLMs’ difficulty is high (0.88) while human annotators’ median announced time is 5 min (low difficulty). Human annotators note that the task is easy, as it is a basic string manipulation with a simple edge case. Observing CodeLLMs’ codes, it is clear that the CodeLLMs struggle with handling the two conditions, especially the last one. Indeed, the CodeLLMs’ code generally messes the two conditions up by merging them (e.g., “Example 1” should be “Example-1”, but the generated code generally returns “Example-__1”).
@@
Listing 8: Example of a task, represented here by a single prompt, difficult for CodeLLMs but judged simple by human-annotators.
Findings 4: Human annotators’ assessments differ from CodeLLMs’ difficulty assessment, emphasising that using human judgement to assess CodeLLMs’ task difficulty might not be a valid proxy. This motivates the usage of approaches like ours that rely directly on CodeLLMs’ assessments to establish characteristics of the tasks.